Find a bond price with f= 100, with seven years of maturity and has a spot yield of 3.6%. Why the answer is $86.81? please solution :)

Calculate the price of a bond with a par value of $1,000 to be paid in 9 years, a coupon rate of 2%, and a required yield of 14.5%. Why the answer is $392.79? please solution :)

If the one-year spot rate is 9.2% and the two-year spot rate is 14% find the forward rate at year 2.. Why the answer is .19.01%? Please, solution :)

Thank you
Answer #1

2.
=Par value*coupon rate/yield*(1-1/(1+yield)^t)+Par value/(1+yield)^t
=1000*2%/14.5%*(1-1/1.145^9)+1000/1.145^9=392.785724790922

3.
=(1+2 year rate)^2/(1+1 year rate)-1
=1.14^2/1.092-1=19.010989010989%

1.
=Face Value/(1+spot yield)^t
=100/1.036^7
=$78.07